OpenStudy (anonymous):

How do I add 1 to the fraction 2/square root of 3

OpenStudy (abb0t):

Start by finding a common denominator.

OpenStudy (anonymous):

\[ 1 + \frac{ 2}{\sqrt 3}= \frac{ 1 \times \sqrt3 +2}{\sqrt 3}= \frac{\sqrt3 +2}{\sqrt 3}\]

OpenStudy (anonymous):

so wouldn't the square roots cancel and just leave me with 2?

OpenStudy (anonymous):

Nope. See here.

OpenStudy (anonymous):